# A chaotic lattice field theory in two dimensions

### v2

@@ I.4 Observables
To understand what this probability is, and motivate the formalism that follows, a bit of quantum-mechanical intuition might be helpful. The semiclassical quantum field theory (for a derivation, see ChaosBook  Appendix A37 WKB quantization [55]) assigns a quantum probability amplitude to a deterministic solution $\mathsf{\Phi}_{c}$[56, 57, 58, 59] with the partition sum
@@
at every lattice site $z$ (see sketch of Fig. 1 (a)). In the WKB approximation, its weight is obtained by expanding the action near the state space point $\mathsf{\Phi}_{c}$ to quadratic order,
@@
II Deterministic field theory
For chaotic (or ‘turbulent’) systems deterministic solutions form a set of state space points, sketched in Fig. 1. In this paper we focus on this set of deterministic solutions. As we show here, despite vastly different appearances, the ‘chaos theory’ of 1970’s is a deterministic Euclidean field theory (see Appendix XII.1).
@@ II Deterministic field theory
II.1 Euclidean field theory
@@
${A}_{\mathbb{A}}[\mathsf{\Phi}]$, the Birkhoff sum Eq. (19) of the observable, or a set of observables, is multiplied by a parameter, or a set of parameters $\beta$. $S[\mathsf{\Phi}]$ is the log probability (in statistics), the Gibbs weight (in statistical physics), or the action (in field theory) of the system under consideration (for examples, see Sec. III).
@@
with field configuration independent bounds $W_{min}[\beta]$, $W_{max}[\beta]$, to be established in Sec. X.1. This requirement is the spatiotemporal generalization of the uniform hyperbolicity of time-evolving dynamical systems, with Lyapunov exponents strictly bounded away from 0.
@@
What we actually need to evaluate is not this integral, but the derivative $W^{\prime}$. That we accomplish in Sec. X.6.
II.2 Deterministic lattice field theory
@@
on every lattice site $z$ (for our examples of defining equations, see Eqs. Eq. (44) – Eq. (46) below). That is all we require, regardless of whether the system has a Lagrangian formulation, or not (for example, Navier-Stokes equations). Deterministic field theory, it turns out, is an elegant, to a novice perhaps impenetrable, definition of what we already know as deterministic chaos and/or turbulence (the precise relation is afforded by Hill’s formulas, derived in Refs. 16, 48).
@@
As any field configuration $\mathsf{\Phi}$ is a point in ${N}_{\mathbb{A}}$-dimensional state space Eq. (12), so is a periodic state $\mathsf{\Phi}_{c}$. Furthermore, just as a temporal evolution period n{n} periodic point is a fixed point of the n{n}th iterate (translation by n{n} temporal lattice sites) of the dynamical time-forward map, every periodic state is a fixed point of a set of symmetries of the theory (Sec. VI.1 and Eq. (152)).
System’s defining equations Eq. (31) are defining equations everyone must obey: look at your left neighbor, right neighbor, remember who you were, make sure you fit in just right. The set $\{\mathsf{\Phi}_{c}\}$ of all possible periodic states is system’s ‘Book of Life’ - a catalogue of all possible ‘lives’, spatiotemporal patterns that defining equations allow, each life a point in system’s state space, each life’s likelihood given by its orbit Jacobian.
@@
In contrast to the quantum action landscape of Fig. 1 (a), for a chaotic (or ‘turbulent’) deterministic system the probability density Fig. 1 (b) is a Dirac porcupine, a set of delta function quills over the primitive cell state space $\mathcal{M}_{\mathbb{A}}$, Eq. (12), one for each solution of defining equations. The primitive cell $\mathbb{A}$ deterministic partition sum Eq. (26) is given by the sum over all periodic states, here labelled ‘$c$’,
@@
is the Birkhoff sum Eq. (19) of observable $a$ over periodic state $\mathsf{\Phi}_{c}$, to be discussed in Sec. X.6. Variants of deterministic partition sums had been computed, by different methods, in different contexts, by many, [62, 60, 63] but always on a given finite primitive cell (geometry), never on the totality of infinite Bravais lattices, as we do here.
@@
evaluated as the ${N}_{\mathbb{A}}\!\times\!{N}_{\mathbb{A}}]$ matrix over the primitive cell $\mathbb{A}$, as the orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A},c}$, to the linear operator Eq. (35), evaluated over infinite Bravais lattice $\mathcal{L}_{\mathbb{A}}$, as the orbit Jacobian operator ${\mathcal{J}}_{c}$ (also called ‘Hessian’, a ‘Jacobi matrix’, or a discrete Schrödinger operator [23, 64, 65]), and to the determinant [46, 47, 56] $\mathrm{Det}\,{\mathcal{J}}_{c}$ in Eq. (33) as the orbit Jacobian (also called ‘discriminant’ or ‘Hill discriminant’ [66]). We add prefix ‘orbit’ to emphasize the distinction between the global orbit stability, and the stability of a forward-in-time evolved state. Our stability exponent $\lambda_{p}$, Eq. (111), is the temporal evolution sum of expanding Floquet exponents generalized to any spacetime dimension.
The orbit Jacobian is the central innovation of our formulation of spatiotemporal chaos, so we discuss at length in sections IV, VIII, IX, X and Appendix B.
II.3 Mosaics
The backbone of a deterministic chaotic system is thus the set of all spatiotemporal solutions of system’s defining equations Eq. (31) that we here refer to as periodic states, or, on occasion, as (multi-)periodic orbits. Depending on the context, in literature they appear under many other names. For example, Gutkin & Osipov [36] refer to a two-dimensional periodic state $\mathsf{\Phi}_{c}$ as a ‘many-particle periodic orbit’, with each lattice site field $\phi_{nt}$ ‘doubly-periodic’, or ‘closed’.
Mosaics.   For a $d$-dimensional spatiotemporal field theory, symbolic description is not a one-dimensional temporal “symbolic dynamics” itinerary, as, for example, a symbol sequence that describes a time-evolving $N$-particle system. The key insight –an insight that applies to coupled-map lattices, and field theories modeled by them, [67, 68, 69, 70, 36, 37] not only systems considered here– is that a field configuration $\mathsf{\Phi}=\{\phi_{z}\}$ over a $d$-dimensional spacetime lattice $z\in\mathbb{Z}^{d}$ is labelled by a finite alphabet symbol lattice $\mathsf{M}=\{m_{z}\}$ over the same $d$-dimensional spacetime lattice.
For field theories studied here, one can partition the values of a lattice site field $\phi_{z}$ into a set of $|\mathcal{A}|$ disjoint intervals, and label each interval by a letter ${m_{z}}\in\mathcal{A}$ drawn from an alphabet $\mathcal{A}$, let’s say
@@ II.3 Mosaics
[…]
@@
elsewhere called ‘symbolic representation’; [67] ‘spatiotemporal code’; [75] ‘symbol tensor’; [76] ‘symbol lattice’; [77, 37] ‘symbol table’; [78] ‘local symbolic dynamics’; [70] ‘symbol block’. [37, 16] A mosaic serves both as a proxy (a ‘name’) for the periodic state $\mathsf{\Phi}_{c}$, and its visualization as color-coded symbol array $\mathsf{M}_{c}$ (for examples, see Fig. 6, Fig. 12 and companion paper III [17]).
@@
If there is only one, distinct mosaic $\mathsf{M}_{c}$ for each periodic state $\mathsf{\Phi}_{c}$, the alphabet is said to be covering. While each periodic state thus gets assigned a unique mosaic that paginates its location in the Book of Life, the converse is in general not true. If a given mosaic $\mathsf{M}$ corresponds to a periodic state, it is admissible, otherwise $\mathsf{M}$ has to be deleted from the list of mosaics. That’s what we do in practice, see Appendix B.1.
@@
In the temporal-evolution setting there is a variety of methods of finding grammar rules that eliminate inadmissible mosaics. While such rules for 2- or higher-dimensional lattice field theories remain, in general, not known to us, we are greatly helped by the observation that in the ‘anti-integrable’ limit [79, 80, 81, 82, 17] (also known as the ‘anti-continuum limit’ in solid state physics, [44] ‘large dissipation limit’ in nonlinear dynamics, [23], ‘weak diffusive coupling’ in stochastic field theory [82]) finite alphabets are known, and offer good starting approximations [17] to the corresponding numerically exact periodic states. For spatiotemporal cat, see Sec. XII.2 Open questions, question  6.
III Examples of spatiotemporal lattice field theories
@@
We shall construct the field theory’s deterministic partition sum (Sec. X) by first enumerating all Bravais lattices (geometries) $\mathcal{L}_{\mathbb{A}}$ (Sec. V), determining prime orbits over each, computing the weight of each (Sec. IX), and then (Sec. X) adding together the contributions of periodic states for each. The potentials may be bounded ($\phi^{4}$ theory) or unbounded ($\phi^{3}$ theory), or the system may be energy conserving or dissipative, as long as the set of its periodic states $\mathsf{\Phi}_{c}$ is bounded in system’s state space Eq. (7). To get a feel for how all this works, we illustrate the theory by applying it to four lattice field theories that we now introduce.
@@
A field theory is defined either by its action, for example a lattice sum over the Lagrangian density for a discretized scalar $d$-dimensional Euclidean $\phi^{k}$ theory, [83, 84, 85, 86, 87, 88, 89]
@@
Defining equations Eq. (31) now take form of second-order difference equations
@@
Here, and in papers I and III [16, 17] we investigate spatiotemporally chaotic lattice field theories using as illustrative examples the $d$-dimensional hypercubic lattice Eq. (6) discretized Klein-Gordon free-field theory, spatiotemporal cat, spatiotemporal $\phi^{3}$ theory, and spatiotemporal $\phi^{4}$ theory, defined respectively by defining equations Eq. (31)
@@
The anti-integrable form [79, 80, 81] of the spatiotemporal $\phi^{3}$, Eq. (45), and spatiotemporal $\phi^{4}$, Eq. (46), is explained in Sec. III.2, and the companion paper III. [17]
The homogeneous, free-field case, Eq. (43) is known as the discretized screened Poisson equation[94, 95] with parameter $\mu$ the reciprocal screening length in the Debye-Hückel or Thomas-Fermi approximations. In statistical mechanics, the related lattice discretized Helmholtz equation is known as the ‘Gaussian model’, [96, 97, 98, 99] and in field theory as the Yukawa or Klein–Gordon equation for a boson of Klein-Gordon (or Yukawa) mass $\mu$. This free-field theory is studied by many, some recent examples are Refs. 100, 63, 101.
@@
III.1 Spatiotemporal cat
The spatiotemporal cat, Eq. (44), that we derive next, is arguably the simplest example of a chaotic (or ‘turbulent’) deterministic field theory, for which all local degrees of freedom are hyperbolic (anti-harmonic, ‘inverted pendula’) rather than oscillatory ‘harmonic oscillators’. We will use it throughout this paper to illustrate our field-theoretic formulation of spatiotemporal chaos.
@@ III.1 Spatiotemporal cat
Theories with lattice site field values compactifed to a circle are known as ‘compact boson’ or ‘compact scalar’ theories, see for example Refs. 104, 105. The ‘mod 1’ in its definition makes the spatiotemporal cat a discontinuous piecewise-linear map, a theory that is nonlinear in the sense that it is not defined globally by a single linear relation, such as the free-field theory Eq. (43), but by a set of distinct piecewise linear conditions, one for each mosaic $\mathsf{M}$.
@@
with ${\mu}^{2}=2({s}-2)$. As in Eq. (41), the orbit Jacobian operator in Eq. (49) can be expressed in terms of shift operators,
@@
In the Hamiltonian, forward-in-time temporal evolution formulation, the dynamics is generated by iterations of a piecewise linear cat map. In the spatiotemporal formulation there is no map, only defining equations, in form of recurrence conditions, so we refer to the three-term recurrence Eq. (50) as the ‘temporal cat’, and to the recurrence condition Eq. (48) in higher spatiotemporal dimensions as the ‘spatiotemporal cat’.
@@
III.2 A many-body example: deterministic $\phi^{4}$ theory

### v3

@@ II.4 Observables
To understand what this probability is, and motivate the formalism that follows, a bit of quantum-mechanical intuition might be helpful. The semiclassical quantum field theory (for a derivation, see Appendix A37 WKB quantization of Ref. 56) assigns a quantum probability amplitude to a deterministic solution $\mathsf{\Phi}_{c}$[57, 58, 59, 60] with the partition sum
@@
at every lattice site $z$ (see the sketch of Fig. 3 (a)). In the WKB approximation, its weight is obtained by expanding the action near the state space point $\mathsf{\Phi}_{c}$ to quadratic order,
@@ III Deterministic field theory
For chaotic (or ‘turbulent’) systems deterministic solutions form a set of state space points, sketched in Fig. 3. In this paper we focus on this set of deterministic solutions. As we show here, despite vastly different appearances, the ‘chaos theory’ of 1970s is a deterministic Euclidean field theory (see Appendix XIII.1).
@@
III.1 Euclidean field theory
@@
${A}_{\mathbb{A}}[\mathsf{\Phi}]$, the Birkhoff sum Eq. (19) of the observable, or a set of observables, is multiplied by a parameter, or a set of parameters $\beta$. $S[\mathsf{\Phi}]$ is the log probability (in statistics), the Gibbs weight (in statistical physics), or the action (in field theory) of the system under consideration (for examples, see Sec. IV).
@@
with field configuration independent bounds $W_{min}[\beta]$, $W_{max}[\beta]$, to be established in Sec. XI.1. This requirement is the spatiotemporal generalization of the uniform hyperbolicity [63] of time-evolving dynamical systems, with Lyapunov exponents strictly bounded away from 0.
@@
What we actually need to evaluate is not this integral, but the derivative $W^{\prime}$. That we accomplish in Sec. XI.6.
III.2 Deterministic lattice field theory
@@
on every lattice site $z$ (for our examples of defining equations, see Eqs. Eq. (44) – Eq. (46)). That is all we require, regardless of whether the system has a Lagrangian formulation, or not (for example, Navier-Stokes equations). Deterministic field theory, it turns out, is an elegant, to a novice perhaps impenetrable, definition of what we already know as deterministic chaos and/or turbulence (the precise relation is afforded by Hill’s formulas, derived in Refs. 16, 48).
@@
As any field configuration $\mathsf{\Phi}$ is a point in ${N}_{\mathbb{A}}$-dimensional state space Eq. (12), so is a periodic state $\mathsf{\Phi}_{c}$. Furthermore, just as a temporal evolution period n{n} periodic point is a fixed point of the n{n}th iterate (translation by n{n} temporal lattice sites) of the dynamical time-forward map, every periodic state is a fixed point of a set of symmetries of the theory (Sec. VII.1 and Eq. (152)).
System’s defining equations Eq. (31) are defining equations everyone must obey: look at your left neighbor, right neighbor, remember who you were, make sure you fit in just right. The set $\{\mathsf{\Phi}_{c}\}$ of all possible periodic states is system’s ‘Book of Life’ - a catalog of all possible ‘lives’, spatiotemporal patterns that defining equations allow, each life a point in system’s state space, each life’s likelihood given by its orbit Jacobian.
@@
In contrast to the quantum action landscape of Fig. 3 (a), for a chaotic (or ‘turbulent’) deterministic system the probability density Fig. 3 (b) is a Dirac porcupine, a set of delta function quills over the primitive cell state space $\mathcal{M}_{\mathbb{A}}$, Eq. (12), one for each solution of defining equations. The primitive cell $\mathbb{A}$ deterministic partition sum Eq. (26) is given by the sum over all periodic states, here labeled ‘$c$’,
@@
is the Birkhoff sum Eq. (19) of observable $a$ over periodic state $\mathsf{\Phi}_{c}$, to be discussed in Sec. XI.6. Variants of deterministic partition sums had been computed, by different methods, in different contexts, by many, [64, 61, 65] but always on a given finite primitive cell (geometry), never on the totality of infinite Bravais lattices, as we do here.
@@
evaluated as the ${N}_{\mathbb{A}}\!\times\!{N}_{\mathbb{A}}]$ matrix over the primitive cell $\mathbb{A}$, as the orbit Jacobian matrix ${\mathcal{J}}_{\mathbb{A},c}$, to the linear operator Eq. (35), evaluated over infinite Bravais lattice $\mathcal{L}_{\mathbb{A}}$, as the orbit Jacobian operator ${\mathcal{J}}_{c}$ (also called ‘Hessian’, a ‘Jacobi matrix’, or a discrete Schrödinger operator [23, 66, 67]), and to the determinant [46, 47, 57] $\mathrm{Det}\,{\mathcal{J}}_{c}$ in Eq. (33) as the orbit Jacobian (also called ‘discriminant’ or ‘Hill discriminant’ [68]). We add prefix ‘orbit’ to emphasize the distinction between the global orbit stability, and the stability of a forward-in-time evolved state. Our stability exponent $\lambda_{p}$, Eq. (111), is the temporal-evolution sum of expanding Floquet exponents generalized to any spacetime dimension.
The orbit Jacobian is the central innovation of our formulation of spatiotemporal chaos, so we discuss at length in sections V, IX, X, XI and Appendix B.
III.3 Mosaics
The backbone of a deterministic chaotic system is, thus, the set of all spatiotemporal solutions of system’s defining equations Eq. (31) that we here refer to as periodic states, or, on occasion, as (multi-)periodic orbits. Depending on the context, in the literature they appear under many other names. For example, Gutkin & Osipov [36] refer to a two-dimensional periodic state $\mathsf{\Phi}_{c}$ as a ‘many-particle periodic orbit’, with each lattice site field $\phi_{nt}$ ‘doubly-periodic’, or ‘closed’.
Mosaics.   For a $d$-dimensional spatiotemporal field theory, symbolic description is not a one-dimensional temporal “symbolic dynamics” itinerary, as, for example, a symbol sequence that describes a time-evolving $N$-particle system. The key insight –an insight that applies to coupled-map lattices, and field theories modeled by them, [69, 70, 71, 72, 36, 37] not only systems considered here– is that a field configuration $\mathsf{\Phi}=\{\phi_{z}\}$ over a $d$-dimensional spacetime lattice $z\in\mathbb{Z}^{d}$ is labeled by a finite alphabet symbol lattice $\mathsf{M}=\{m_{z}\}$ over the same $d$-dimensional spacetime lattice.
For field theories studied here, one can partition the values of a lattice site field $\phi_{z}$ into a set of $|\mathcal{A}|$ disjoint intervals, and label each interval by a letter ${m_{z}}\in\mathcal{A}$ drawn from an alphabet $\mathcal{A}$, let us say
@@ III.3 Mosaics
This associates a $d$-dimensional mosaic $\mathsf{M}_{c}$ to a field configuration $\mathsf{\Phi}_{c}$ over $d$-dimensional lattice [73, 74, 75, 76]
@@
elsewhere called ‘symbolic representation’; [69] ‘spatiotemporal code’; [77] ‘symbol tensor’; [78] ‘symbol lattice’; [79, 37] ‘symbol table’; [80] ‘local symbolic dynamics’; [72] ‘symbol block’. [37, 16] A mosaic serves both as a proxy (a ‘name’) for the periodic state $\mathsf{\Phi}_{c}$, and its visualization as color-coded symbol array $\mathsf{M}_{c}$ (for examples, see Fig. 4, Fig. 5, companion paper III, [17] and Lego mosaics of Ref. 81).
Figure 4: Examples of spatiotemporal mosaic tilings Eq. (37) of $6\!\times\!6]_{0}$} primitive cell by repeats of smaller prime periodic states. (a) $3\!\times\!1]_{0}$} temporally steady state. (b) $1\!\times\!3]_{0}$} spatially steady state. (c) $2\!\times\!1]_{1}$} relative-periodic prime orbit, spatial period-2, temporal period-2; compare with Fig. 6 (a). (d) $3\!\times\!1]_{1}$} relative-periodic prime orbit, spatial period-3, temporal period-3. (e) $3\!\times\!2]_{0}$} spatial period-3, temporal period-2. (f) $3\!\times\!2]_{1}$} of Figs. 6 (b) and 8. It is a relative-periodic prime orbit, of spatial period-3, temporal period-6. See also Fig. 5 and Appendix A.
(a)    Refer to caption
(b)    Refer to caption
Figure 5: Mosaics Eq. (37) of two $18\!\times\!18]_{0}$ spatiotemporal cat periodic states which share the sub-mosaic within the $12\!\times\!12]$ region enclosed by the black square, and have different symbols outside the sub-mosaic. Color-coded 8-letter alphabet Eq. (163), $\mu^{2}=1$. This will be discussed further in Fig. 13.
@@
If there is only one, distinct mosaic $\mathsf{M}_{c}$ for each periodic state $\mathsf{\Phi}_{c}$, the alphabet is said to be covering. While each periodic state thus gets assigned a unique mosaic that paginates its location in the Book of Life, the converse is in general not true. If a given mosaic $\mathsf{M}$ corresponds to a periodic state, it is admissible; otherwise $\mathsf{M}$ has to be deleted from the list of mosaics. That’s what we do in practice, see Appendix B.1.
@@
In the temporal-evolution setting there is a variety of methods of finding grammar rules that eliminate inadmissible mosaics. While such rules for 2- or higher-dimensional lattice field theories remain, in general, not known to us, we are greatly helped by the observation that in the ‘anti-integrable’ limit [82, 83, 84, 85, 17] (also known as the ‘anti-continuum limit’ in solid state physics, [44] ‘large dissipation limit’ in nonlinear dynamics, [23], ‘weak diffusive coupling’ in stochastic field theory [85]) finite alphabets are known, and offer good starting approximations [17] to the corresponding numerically exact periodic states. For spatiotemporal cat, see Sec. XIII.2, question 6.
IV Examples of spatiotemporal lattice field theories
@@
We shall construct the field theory’s deterministic partition sum (Sec. XI) by first enumerating all Bravais lattices (geometries) $\mathcal{L}_{\mathbb{A}}$ (Sec. VI), determining prime orbits over each, computing the weight of each (Sec. X), and then (Sec. XI) adding together the contributions of periodic states for each. The potentials may be bounded ($\phi^{4}$ theory) or unbounded ($\phi^{3}$ theory), or the system may be energy conserving or dissipative, as long as the set of its periodic states $\mathsf{\Phi}_{c}$ is bounded in system’s state space Eq. (7). To get a feel for how all this works, we illustrate the theory by applying it to four lattice field theories that we now introduce.
@@
A field theory is defined either by its action, for example a lattice sum over the Lagrangian density for a discretized scalar $d$-dimensional Euclidean $\phi^{k}$ theory, [86, 87, 88, 89, 90, 91, 92, 93]
@@
Defining equations Eq. (31) now take the form of second-order difference equations
@@
Here, and in papers I and III [16, 17] we investigate spatiotemporally chaotic lattice field theories using as illustrative examples the $d$-dimensional hypercubic lattice Eq. (6) discretized Klein-Gordon free-field theory, spatiotemporal cat, spatiotemporal $\phi^{3}$ theory, and spatiotemporal $\phi^{4}$ theory, defined, respectively by defining equations Eq. (31)
@@
The anti-integrable form [82, 83, 84] of the spatiotemporal $\phi^{3}$, Eq. (45), and spatiotemporal $\phi^{4}$, Eq. (46), is explained in Sec. IV.2, and the companion paper III. [17]
The homogeneous, free-field case, Eq. (43) is known as the discretized screened Poisson equation[98, 99] with parameter $\mu$ the reciprocal screening length in the Debye-Hückel or Thomas-Fermi approximations. In statistical mechanics, the related lattice discretized Helmholtz equation is known as the ‘Gaussian model’, [100, 101, 102, 103] and in field theory as the Yukawa or Klein–Gordon equation [104] for a boson of Klein-Gordon (or Yukawa) mass $\mu$. This free-field theory is studied by many, some recent examples are Refs. 105, 65, 106.
@@
IV.1 Spatiotemporal cat
The spatiotemporal cat, Eq. (44), that we derive next, is arguably the simplest example of a chaotic (or ‘turbulent’) deterministic field theory, for which all local degrees of freedom are hyperbolic (anti-harmonic, ‘inverted pendula’) rather than oscillatory ‘harmonic oscillators’. [107] We will use it throughout this paper to illustrate our field-theoretic formulation of spatiotemporal chaos.
@@ IV.1 Spatiotemporal cat
Theories with lattice site field values compactified to a circle are known as ‘compact boson’ or ‘compact scalar’ theories, see for example Refs. 110, 111. The ‘mod 1’ in its definition makes the spatiotemporal cat a discontinuous piecewise-linear map, a theory that is nonlinear in the sense that it is not defined globally by a single linear relation, such as the free-field theory Eq. (43), but by a set of distinct piecewise linear conditions, one for each mosaic $\mathsf{M}$.
@@
with ${\mu}^{2}=2({s}-2)$[107] As in Eq. (41), the orbit Jacobian operator in Eq. (49) can be expressed in terms of shift operators,
@@
In the Hamiltonian, forward-in-time temporal evolution formulation, the dynamics is generated by iterations of a piecewise linear cat map. In the spatiotemporal formulation there is no map, only defining equations, in the form of recurrence conditions, so we refer to the three-term recurrence Eq. (50) as the temporal cat, and to the recurrence condition Eq. (48) in higher spatiotemporal dimensions as the spatiotemporal cat.
@@
IV.2 A many-body example: deterministic $\phi^{4}$ theory